# ISM: individual objects (N132D) — ISM: supernova remnants — X-rays: ISM The XRISM First Light Observation: Velocity Structure and Thermal Properties of the Supernova Remnant N132D

### v1

@@ Abstract
We present an initial analysis of the XRISM first-light observation of the supernova remnant (SNR) N132D in the Large Magellanic Cloud. The Resolve microcalorimeter has obtained the first high-resolution spectrum in the 1.6–10 keV band, which contains K-shell emission lines of Si, S, Ar, Ca, and Fe. We find that the Si and S lines are relatively narrow, with a broadening represented by a Gaussian-like velocity dispersion of km s-1. The Fe He lines are, on the other hand, substantially broadened with km s-1. This broadening can be explained by a combination of the thermal Doppler effect due to the high ion temperature and the kinematic Doppler effect due to the SNR expansion. Assuming that the Fe He emission originates predominantly from the supernova ejecta, we estimate the reverse shock velocity at the time when the bulk of the Fe ejecta were shock heated to be (in the observer frame). We also find that Fe Ly emission is redshifted with a bulk velocity of km s-1, substantially larger than the radial velocity of the local interstellar medium surrounding N132D. These results demonstrate that high-resolution X-ray spectroscopy is capable of providing constraints on the evolutionary stage, geometry, and velocity distribution of SNRs.
@@ 1 Introduction
The SNR N132D, located in the bar of the Large Magellanic Cloud (LMC), is one of the most studied objects of its class at virtually every wavelength: e.g., radio (Dickel & Milne, 1995; Sano et al., 2020), infrared (Williams et al., 2006; Rho et al., 2023), and gamma-ray (Acero et al., 2016; Ackermann et al., 2016; H.E.S.S. Collaboration, 2021). Its periphery has an elliptical shape of along the major and minor axes, respectively, which corresponds to 27 pc 19 pc at the distance to the LMC of 50 kpc (Pietrzyński et al., 2013). N132D is thought to be in transitional phase from young to middle-aged (e.g., Bamba et al. (2018)). Therefore, both swept-up ISM and ejecta contribute to its electromagnetic spectrum.
@@ 2.2 Xtend
The Xtend instrument was operating in the full-window mode during the first observation and in the 1/8-window mode in the second observation, partly for calibration purposes. After the standard event screening, we obtained effective exposures of 123 ks and 71 ks for the first and second observations, respectively. At the time of the observations of N132D, one of the charge-injection rows overlapped with the aim point when the instrument was operating in the full-window mode (Figure 1a)444This issue was fixed in February 2024, so the charge-injection row is no longer overlapping with the aim point.. Therefore, we use only the second observation data for imaging analysis. Figure 1b shows a three-color image of N132D obtained from the second observation: red, green, and blue correspond to 0.3–0.5 keV, 0.5–1.75 keV, and 1.75–10 keV, respectively.
@@ 3 Analysis
Figure 2 shows the Resolve spectrum in the 1.6–10 keV band extracted from the entire field of view of arcmin2 or 35 pixels. The K-shell emission lines of Si, S, Ar, Ca, and Fe are detected with high significance. The presence of the Fe Ly emission, first suggested in the Suzaku study (Bamba et al., 2018), is also confirmed. A remarkable difference is found between the He-like emission lines of the intermediate mass elements (IMEs) and Fe. The former are relatively narrow so the forbidden and resonance lines are well resolved, whereas the latter is significantly broadened. This is not an instrumental effect, since the spectral resolution (measured in eV) is approximately constant across the 1.6–10 keV band. It is therefore indicative that the IME and Fe emission originate from different plasma components, consistent with the Hitomi study.
@@ 3.3 Fullband spectrum
∗*∗∗*∗footnotemark: The abundances are linked among the different temperature components.
@@ 3.4 Narrow band image
The result based on Model B in the previous subsection implies that the Fe He and Ly emission originates from different plasma components. If this hypothesis is true, different spatial distributions of these emissions should be expected. We thus generate narrow band images of the Fe He and Ly emission as well as the S He emission to search for morpological differences. The results are shown in Figure 6; the top and bottom panels are the Resolve and Xtend images, respectively. The distribution of the Fe Ly emission is localized near the SNR center, whereas the Fe He emission is more widely distributed. This difference is seen by both instruments, although the low statistics in the emission lines makes it difficult to determine their true morphology. The morphological difference between the S and Fe emission is also confirmed, consistent with previous studies (Behar et al., 2001; Borkowski et al., 2007; Sharda et al., 2020).
@@ 4.1 Radial velocity
The large redshift observed in the Fe Ly emission implies that this emission originates from the Fe-rich ejecta with a highly asymmetric velocity distribution. The ejecta scenario is also supported from the high electron temperature that can be achieved by a high-velocity shock. The morphology of the Fe Ly emission is centrally concentrated (Figure 6), suggesting that this hot Fe ejecta component is present only at the far side of the SNR. Theoretically, an ionization timescale of cm-3 s is required to produce a sufficient fraction of H-like Fe in an ionizing plasma. Therefore, from the estimated SNR age of 2770 yr (Banovetz et al., 2023), the electron density of this component is estimated to be cm-3. Such a density implies that the Fe ejecta form a compact knot. In fact, the normalization of the very-high- component given in Table 3 corresponds to the emitting volume of cm3, less than 0.1% of the total SNR volume ( cm3).
@@ 5 Conclusions
We have presented analysis of the XRISM first-light observation data of N132D, the X-ray brighest SNR in the LMC. The excellent performance of the X-ray microcalorimeter Resolve has enabled us to perform high resolution spectroscopy of this SNR in the 1.6–10-keV band, for the first time. Our analysis has revealed that the K-shell emission lines of Si and S, whose origin is thought to be the swept-up ISM, are mildly broadened with km s-1. Under an assumption of a nearly symmetrically expanding ISM shell, the radial component of the blast wave velocity is estimated to be km s-1, lower than the recent proper motion measurement with Chandra. On the other hand, the Fe H complex emission is substantially broadened with km s-1. If this emission originates from the ejecta, the observed line width can be explained through a combination of the thermal Doppler broadening due to the high ion temperature in the non-equilibrium plasma and kinematic Doppler effect due to the expansion of the shocked ejecta. We have also provided an estimate for the evolutionary stage of the SNR, putting a constraint on the reverse shock velocity (in observer frame) to be , the value at the time when the bulk of the Fe ejecta were shock heated. The redshift observed in the Si, S, and Fe He lines is equivalent to the radial velocity of the ISM surrounding N132D ( km s-1), whereas that observed in Fe Ly indicates a substantially larger radial velocity of 890 km s-1, about 600 km s-1 greater than the radial velocity of the local ISM in the LMC. Also our imaging analysis revealed that the Fe Ly emission is concentrated around the SNR center. These results suggest that the Fe Ly emission originates from hot Fe ejecta that are present only at the far side of the SNR, distinct from the component responsible for the Fe He emission.
@@
This work was supported by JSPS KAKENHI grant numbers JP22H00158, JP22H01268, JP22K03624, JP23H04899, JP21K13963, JP24K00638, JP24K17105, JP21K13958, JP21H01095, JP23K20850, JP24H00253, JP21K03615, JP24K00677, JP20K14491, JP23H00151, JP19K21884, JP20H01947, JP20KK0071, JP23K20239, JP24K00672, JP24K17104, JP24K17093, JP20K04009, JP21H04493, JP20H01946, JP23K13154, JP19K14762, JP20H05857, and JP23K03459, and NASA grant numbers 80NSSC20K0733, 80NSSC18K0978, 80NSSC20K0883, 80NSSC20K0737, 80NSSC24K0678, 80NSSC18K1684, and 80NNSC22K1922. LC acknowledges support from NSF award 2205918. CD acknowledges support from STFC through grant ST/T000244/1. LG acknowledges financial support from Canadian Space Agency grant 18XARMSTMA. AT and the present research are in part supported by the Kagoshima University postdoctoral research program (KU-DREAM). SY acknowledges support by the RIKEN SPDR Program. IZ acknowledges partial support from the Alfred P. Sloan Foundation through the Sloan Research Fellowship. Part of this work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. The material is based upon work supported by NASA under award number 80GSFC21M0002. This work was supported by the JSPS Core-to-Core Program, JPJSCCA20220002. The material is based on work supported by the Strategic Research Center of Saitama University.

### v2

@@ Abstract
We present an initial analysis of the X-Ray Imaging and Spectroscopy Mission (XRISM) first-light observation of the supernova remnant (SNR) N132D in the Large Magellanic Cloud. The Resolve microcalorimeter has obtained the first high-resolution spectrum in the 1.6–10 keV band, which contains K-shell emission lines of Si, S, Ar, Ca, and Fe. We find that the Si and S lines are relatively narrow, with a broadening represented by a Gaussian-like velocity dispersion of km s-1. The Fe He lines are, on the other hand, substantially broadened with km s-1. This broadening can be explained by a combination of the thermal Doppler effect due to the high ion temperature and the kinematic Doppler effect due to the SNR expansion. Assuming that the Fe He emission originates predominantly from the supernova ejecta, we estimate the reverse shock velocity at the time when the bulk of the Fe ejecta were shock heated to be (in the observer frame). We also find that Fe Ly emission is redshifted with a bulk velocity of km s-1, substantially larger than the radial velocity of the local interstellar medium surrounding N132D. These results demonstrate that high-resolution X-ray spectroscopy is capable of providing constraints on the evolutionary stage, geometry, and velocity distribution of SNRs.
@@ 1 Introduction
The SNR N132D, located in the bar of the Large Magellanic Cloud (LMC), is one of the most studied objects of its class at virtually every wavelength: e.g., radio (Dickel & Milne, 1995; Sano et al., 2020), infrared (Williams et al., 2006; Rho et al., 2023), and gamma-ray (Acero et al., 2016; Ackermann et al., 2016; H.E.S.S. Collaboration, 2021). Its periphery has an elliptical shape of along the major and minor axes, respectively, which corresponds to 27 pc 19 pc at the distance to the LMC of 50 kpc (Pietrzyński et al., 2013). N132D is thought to be in a transitional phase from young to middle-aged (e.g., Bamba et al. (2018)). Therefore, both swept-up ISM and ejecta contribute to its electromagnetic spectrum.
@@ 2.2 Xtend
The Xtend instrument was operating in the full-window mode during the first observation and in the 1/8-window mode in the second observation, partly for calibration purposes. After the standard event screening, we obtained effective exposures of 123 ks and 71 ks for the first and second observations, respectively. At the time of the observations of N132D, one of the charge-injection rows overlapped with the aim point when the instrument was operating in the full-window mode (Figure 1a)444This issue was fixed in 2024 March, so the charge-injection row is no longer overlapping with the aim point.. Therefore, we use only the second observation data for imaging analysis. Figure 1b shows a three-color image of N132D obtained from the second observation: red, green, and blue correspond to 0.3–0.5 keV, 0.5–1.75 keV, and 1.75–10 keV, respectively.
@@ 3 Analysis
Figure 2 shows the Resolve spectrum in the 1.6–10 keV band extracted from the entire field of view of arcmin2 or 35 pixels. The K-shell emission lines of Si, S, Ar, Ca, and Fe are detected with high significance. The presence of the Fe Ly emission, first suggested in the Suzaku study (Bamba et al., 2018), is also confirmed. A remarkable difference is found between the He-like emission lines of the intermediate mass elements (IMEs) and Fe. The former are relatively narrow so the forbidden and resonance lines are well resolved, whereas the latter is significantly broadened. This is not an instrumental effect, since the spectral resolution (measured in eV) is approximately constant across the 1.6–10 keV band. It is therefore indicative that the IME and Fe emission originate from different plasma components.
@@ 3.3 Fullband spectrum
∗*∗∗*∗footnotemark: The abundances are linked between the different temperature components.
@@ 3.4 Narrow band image
The result based on Model B in the previous subsection implies that the Fe He and Ly emissions originate from different plasma components. If this hypothesis is true, different spatial distributions of these emissions should be expected. We thus generate narrow band images of the Fe He and Ly emission as well as the S He emission to search for morpological differences. The results are shown in Figure 6; the top and bottom panels are the Resolve and Xtend images, respectively. The distribution of the Fe Ly emission is localized near the SNR center, whereas the Fe He emission is more widely distributed. This difference is seen by both instruments, although the low statistics in the emission lines makes it difficult to determine their true morphology. The morphological difference between the S and Fe emission is also confirmed, consistent with previous studies (Behar et al., 2001; Borkowski et al., 2007; Sharda et al., 2020).
@@ 4.1 Radial velocity
The large redshift observed in the Fe Ly emission implies that this emission originates from the Fe-rich ejecta with a highly asymmetric velocity distribution. The ejecta scenario is also supported from the high electron temperature that can be achieved by a high-velocity shock. The morphology of the Fe Ly emission is centrally concentrated (Figure 6), suggesting that this hot Fe ejecta component is present only at the far side of the SNR. Theoretically, an ionization timescale of cm-3 s is required to produce a sufficient fraction of H-like Fe in an ionizing plasma. Therefore, from the estimated SNR age of 2770 yr (Banovetz et al., 2023), the electron density of this component is estimated to be cm-3. Such a high density implies that the Fe ejecta form a compact knot. In fact, the normalization of the very-high- component given in Table 3 corresponds to the emitting volume of cm3, less than 0.1% of the total SNR volume ( cm3).
@@ 5 Conclusions
We have presented analysis of the XRISM first-light observation data of N132D, the X-ray brighest SNR in the LMC. The excellent performance of the X-ray microcalorimeter Resolve has enabled us to perform high resolution spectroscopy of this SNR in the 1.6–10-keV band, for the first time. Our analysis has revealed that the K-shell emission lines of Si and S, whose origin is thought to be the swept-up ISM, are mildly broadened with km s-1. Under an assumption of a nearly symmetrically expanding ISM shell, the radial component of the blast wave velocity is estimated to be km s-1, lower than the recent proper motion measurement with Chandra. On the other hand, the Fe H complex emission is substantially broadened with km s-1. If this emission originates from the ejecta, the observed line width can be explained through a combination of the thermal Doppler broadening due to the high ion temperature in the non-equilibrium plasma and kinematic Doppler effect due to the expansion of the shocked ejecta. We have also provided an estimate for the evolutionary stage of the SNR, putting a constraint on the reverse shock velocity (in observer frame) to be , the value at the time when the bulk of the Fe ejecta were shock-heated. If km s-1, which is reasonably expected from the theoretical point of view, the current Fe temperature is estimated to be MeV. The redshift observed in the Si, S, and Fe He lines is equivalent to the radial velocity of the ISM surrounding N132D ( km s-1), whereas that observed in Fe Ly indicates a substantially larger radial velocity of 890 km s-1, about 600 km s-1 greater than the radial velocity of the local ISM in the LMC. Also our imaging analysis revealed that the Fe Ly emission is concentrated around the SNR center. These results suggest that the Fe Ly emission originates from hot Fe ejecta that are present only at the far side of the SNR, distinct from the component responsible for the Fe He emission.
@@
This work was supported by JSPS KAKENHI grant numbers JP22H00158, JP22H01268, JP22K03624, JP23H04899, JP21K13963, JP24K00638, JP24K17105, JP21K13958, JP21H01095, JP23K20850, JP24H00253, JP21K03615, JP24K00677, JP20K14491, JP23H00151, JP19K21884, JP20H01947, JP20KK0071, JP23K20239, JP24K00672, JP24K17104, JP24K17093, JP20K04009, JP21H04493, JP20H01946, JP23K13154, JP19K14762, JP20H05857, JP23K03459, and JP22KJ1047, and NASA grant numbers 80NSSC20K0733, 80NSSC18K0978, 80NSSC20K0883, 80NSSC20K0737, 80NSSC24K0678, 80NSSC18K1684, and 80NNSC22K1922. LC acknowledges support from NSF award 2205918. CD acknowledges support from STFC through grant ST/T000244/1. LG acknowledges financial support from Canadian Space Agency grant 18XARMSTMA. AT and the present research are in part supported by the Kagoshima University postdoctoral research program (KU-DREAM). SY acknowledges support by the RIKEN SPDR Program. IZ acknowledges partial support from the Alfred P. Sloan Foundation through the Sloan Research Fellowship. Part of this work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. The material is based upon work supported by NASA under award number 80GSFC21M0002. This work was supported by the JSPS Core-to-Core Program, JPJSCCA20220002. The material is based on work supported by the Strategic Research Center of Saitama University.